OpenStudy (peterg2001):

help please will medal Which of the following points is a solution of y ≤ -|x| - 1? (-1, -3) (0, 0) (1, -1)

OpenStudy (peterg2001):

@sparrow2 @wwhitlock

OpenStudy (peterg2001):

@Mahmood786123

OpenStudy (peterg2001):

c?

OpenStudy (dayakar):

substitute x , y values in the inequality and check for which values it is true for example 1) (x,y) = (-1 ,-3) \[y \le - \left| x \right|-1\] \[-3 \le -\left| -1 \right|-1\] \[-3 \le -1 -1\] \[-3 \le -2\] is true or false

OpenStudy (peterg2001):

i dont know

OpenStudy (dayakar):

like above check each one u can get the answer